OpenStudy (anonymous):

if a sphere has a radius of 6 and a surface area of 144 pi inches^2. if the radius is doubled what will happen to the surface area?

OpenStudy (shane_b):

The surface area of a sphere is given by the formula:\[SA_{sphere}=4\pi r^2\]So what happens if you double r?

OpenStudy (anonymous):

just 144*4=576inches^2

OpenStudy (shane_b):

No...it won't be doubled...it will be quadrupled

OpenStudy (shane_b):

If you double any squared value you quadruple the value.

OpenStudy (anonymous):

Shane is right... (2r)^2=4r^2 The surface area would be multiplied by 4.